Mehanical engineering, Kinematics and dynamics


It is asked to model a general scheme of 3 machines in series and to express the formula of global efficiency. Moreover a scheme is asked to sketch a system with three gear transmissions in series and to calculate the global efficiency when considering that the three gear transmission have efficiency. equal to 0.7, 0.8 , 0.9 respectively. Finally the output power distribution is calculated from an input of 100MW.
Answer #1

Given:
Three gear transmissions in series with efficiencies 0.7,0.8 and 0.9.
Now,
The global efficiency will be product of all efficiencies for series combination.
Therefore,
Global efficiency=0.7×0.8×0.9=0.504

Power output=Global efficiency×Power input= 0.504×100= 50.4MW